Ticket Number: HEA-NOC/20020115-01 Ticket Status: UPDATE
Ticket Type: Unscheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: HEAnet <==> UCD
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 20020115 04:11 UTC
Ticket Closed: Problem End: 20020115 09:50 UTC
Routing flaps on UCD access circuit (AVC500059 on bearer ATM502601)
with loss of protocol and IP connectivity.
All traffic to and from UCD. Mail backed up by TCD and Esat (equal
20020115 07:30 Report fault to Eircom.
200201515 08:20 Eircom report anomalous output from UCD on ceased VC.
20020115 09:25 UCD investigating problem at their border router.
20020115 09:50 UCD bring up HEAnet link and are investigating
cause of outage.
20020115 14:00 UCD report identification of source and resolution
of service outage; monitoring continues.
Time to Fix:
5 hours 39 minutes